In developing a Bar X chart, the centerline of the Bar X chart is determined to...

  1. In developing a Bar X chart, the centerline of the Bar X chart is determined to be 0.4949, the centerline of the R chart is determined to be 0.0130, and the factor A2 used in the formulas to calculate the lower and upper limits of the Bar X chart is determined to be 0.7290. Based on these information,

    a.

    The lower control limit of the Bar X chart is 0.4854

    b.

    The upper control limit of the Bar X chart is 0.5044

    c.

    a and b

    d.

    Not enough information given to answer this question
Answer #1

Ceisen, R = 0.0130 F= = 0.4949 , Az 0.729 Now, UCL X+Az Ř = 0.4949+0.729 (0.013) = 0.5044 LCC; = 7 - AzR = 0.4949 – 0.729 (0.